OpenStudy (anonymous):

Solve. 1 over 25 = 5x + 4

OpenStudy (anonymous):

Solve. 1 over 25 = 5^x + 4

OpenStudy (anonymous):

1/4.2 or do you need a decimal?

OpenStudy (anonymous):

UHH these are the options, x = –6 x = negative 7 over 2 x = 9 over 2 x = 2

OpenStudy (anonymous):

Im so sorry, I have to go I cant help you but i'll be able to eliminate the last choice for you :D BYe

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

1/25 = 5^-2 So, applying logarithms on both sides, you get x+4 = -2. So, x = -6.

OpenStudy (anonymous):

thanks so much